OpenStudy (anonymous):

Select the fraction that is equivalent to 42 .8181

OpenStudy (anonymous):

471/11

hero (hero):

x = .8181... 100x = 81.8181... 100x - x = 81.8181... - .8181 99x = 81 x = 81/99 Therefore 42.8181 = 42 and 81/99 or simply \[42\frac{81}{99}\]